OpenStudy (animelover):

Will Fan Metal! Drag the steps to place them in the correct order to solve the equation. 13−0.5(2x−6)+2x=17 13+3−x+2x=17 x=1 13−0.5⋅2x−0.5⋅(−6)+2x=17 16+x=17 13−x+3+2x=17

OpenStudy (animelover):

13−0.5(2x−6)+2x=1713−0.5⋅2x−0.5⋅(−6)+2x=1713−x+3+2x=1713+3−x+2x=1716+x=17x=1